OpenStudy (anonymous):

find f-g given that: f(x)=x^3-3x+2√x-7, g(x)=-6x^3+10x-√x

OpenStudy (anonymous):

put a minus sign between them use parentheses

OpenStudy (anonymous):

\[x^3-x+2\sqrt x-7-(-6x^3+10x-\sqrt x)\] is a start

OpenStudy (anonymous):

then distribute the minus sign \[x^3-3x+2\sqrt x-7+6x^3-10x+\sqrt x\] and finally combine like terms

OpenStudy (anonymous):

is that the answer ? because I don't see that as one of the choices

OpenStudy (anonymous):

7x^3-13x+3√x-7